# Robust Global Fréchet Regression via Weight Regularization

### v1

@@ Abstract
To assess the performance in the distribution response. For each observation $i=1,\ldots,n$, the covariate $X_{i}$ is independently drawn from a uniform distribution on $[0,1]$. Conditional on $X_{i}$, the true mean parameter $\mu_{i}$ is generated from the normal distribution, $\mu_{i}|X_{i}\sim\mathcal{N}(\mu_{0}+\beta X_{i},v_{1})$. Also, the true standard deviation parameter $\sigma$ is generated from a gamma distribution, $\sigma_{i}|X_{i}\sim{\rm Ga}(\alpha_{i},\lambda_{i}$, where $\alpha_{i}=(\sigma_{0}+\gamma X_{i})^{2}/v_{2}$ is a shape parameter and $\lambda_{i}=v_{2}/(\sigma_{0}+\gamma X_{i})$ is a rate parameter. For a fixed set of quantile levels $\{z_{j}\}$ as an equally spaced sequence starting from 0.1 to 0.9 with an increment of 0.01, expressed as $z_{j}=0.1+0.01\times(j-1)$ for $j=1,2,\ldots,81$. The response variable $Y_{ij}$ at quantile $z_{j}$ is generated via the quantile function of the standard normal distribution as $Y_{ij}=\mu_{i}+\sigma_{i}\Phi^{-1}(z_{j})$, where $\Phi(\cdot)$ denotes the standard normal distribution function. To account for contaminated data, we randomly selected a predetermined number of samples from the observations and introduced a constant Shift value to the corresponding response variable at each. Regarding generating contaminated data, we adopted the same strategy as with the matrix response. We randomly select a 10% or 20% dataset and add a 50 or 100 Shift in each element of the corresponding distribution observation.
@@
Table 4 summarizes average MISE values averaged over 100 Monte Carlo replications with Monte Carlo standard errors (in parentheses) for the original and robust Fréchet regressions under distribution responses with $n=50$ and $n=100$ across two contamination scenarios. In the absence of contamination, both methods yield comparable MISEs. As the contamination proportion increases, MISEs rise for both methods, but the increase is markedly more pronounced for the original Fréchet regression, especially under large contamination and shifts. For example, at $n=50$ with a contamination proportion of $0.2$ and shift of $100$, the MISE of the robust Fréchet regression is approximately a quarter of that of the original Fréchet regression. Similar to $n=100$, underscoring the superior stability of the robust Fréchet regression in the presence of outliers. The $\lambda$ and $\gamma$ results for the distribution response are also reported in Table 5). Consistent with the matrix-response case, $\gamma$ and $\lambda$ remain large when greater shift value and higher contamination proportions.
@@
Table 5: The value of $\lambda$ and $\gamma$ for distribution response under various data configurations (sample size $n$ and other parameter settings). Values are mu;multiplied by $10^{4}$.
@@ Discussion
However, our method still has certain limitations. Specifically, it only assesses the overall outlierness for each observation as a whole. It does not allow for the evaluation of the outlierness of individual components within each observation. For instance, in the case of a matrix response, the anomaly of a single component may lead to the entire matrix being identified as an outlier. Nevertheless, in the estimation process, the presence of other normal components can help decrease the variance of the model. Therefore, extensions addressing this limitation will be considered in our future work.
@@ Appendix A Proof of Proposition 1
under $W_{i}\geq 0$, and the above objective function is strictly convex. When $1-W_{i}>0$, the objective function can be expressed as $\gamma(W_{i}-\bar{w}_{i})^{2}-\gamma\bar{w}_{i}^{2}+\lambda+\gamma$, where
$\bar{w}_{i}=1-\frac{r_{i}(u)-\lambda}{2\gamma}.$
When $0<\bar{w}_{i}<1$, namely, $\lambda<r_{i}(u)\leq\lambda+2\gamma$, $\bar{w}_{i}$ itself is the optimal value of $W_{i}$. Moreover, the optimal $W_{i}$ is 0 when $\bar{w}_{i}\leq 0$, namely, $r_{i}(u)\geq\lambda+2\gamma$, and $1$ when $\bar{w}_{i}\geq 1$, namely, $r_{i}(u)\leq\lambda$. Also, when $1-W_{i}\leq 0$, the objective function is $\{r_{i}(u)+\lambda\}W_{i}+\gamma(1-W_{i})^{2}$ as a function of $W_{i}$, which is increasing on $W_{i}\geq 1$ and the minimizer is ${\widetilde{W}}_{i}(u)=1$.
@@ Appendix B Proof of Proposition 2
$|\widetilde{W}_{i}(u_{1})-\widetilde{W}_{i}(u_{2})|\leq\frac{1}{2\gamma}|r_{i}(u_{1})-r_{i}(u_{2})|\leq\frac{g(X_{i},x)L_{d}}{2\gamma}d(u_{1},u_{2}).$
Hence, it holds that $\|\widetilde{w}(u_{1})-\widetilde{w}(u_{2})\|\leq(2\gamma)^{-1}D_{g}L_{d}d(u_{1},u_{2})$. Under (C3), it holds that
@@
where $\rho=C_{u}D_{g}L_{d}/2\gamma$. When $\rho<1$, $T$ is a contraction mapping, which completes the proof.
Appendix C Regularity Conditions for Specific Models
@@
$\|A(w_{1})-A(w_{2})\|_{F}$ $\leq\|w_{1}-w_{2}\|\Big\{\sum_{i=1}^{n}\|g(X_{i},x)Y_{i}\|_{F}^{2}\Big\}^{1/2}\leq\sqrt{n}\|w_{1}-w_{2}\|D_{g,\infty}D_{u},$
where $D_{g,\infty}=\max_{i=1,\ldots,n}|g(X_{i},x)|$. We also note that
$\|A(w_{2})\|_{F}\leq\sum_{i=1}^{n}W_{i}|g(X_{i},x)|\cdot\|Y_{i}\|_{F}\leq nD_{g,\infty}D_{u},$
and $|S(w_{2})-S(w_{1})|\leq\sqrt{n}D_{g,\infty}\|w_{2}-w_{1}\|$ from the Cauchy-Schwartz inequality. Then, we have
$\|\Phi(w_{1})-\Phi(w_{2})\|_{F}\leq\left(\frac{n^{\frac{1}{2}}D_{g,\infty}D_{u}}{S_{\min}}+\frac{n^{\frac{3}{2}}D^{2}_{g,\infty}D_{u}}{S_{\min}^{2}}\right)\|w_{2}-w_{1}\|,$
@@
According to the definition of $L_{2}$-Wasserstein distance. we obtain
$|d^{2}(Y_{i},u_{1})-d^{2}(Y_{i},u_{2})|$ $=\bigl(\|u-F^{-1}_{i}(z)\|_{2}-\|u-F^{-1}_{i}(z)\|_{2}\bigr)\bigl(\|u-F^{-1}_{i}(z)\|_{2}+\|u-F^{-1}_{i}(z)\|_{2}\bigr)$
$\leq 2D_{W}\,d(u_{1},u_{2})$
similar as defined in Proofing Matrix response whereby (C3) is satisfied with $L_{d}=2D_{W}$. Furthermore, for the distribution response, we also define: $A(w)=\sum_{i=1}^{n}W_{i}g(X_{i},x)F^{-1}_{i}(z)$ and $S(w)=\sum_{i=1}^{n}W_{i}g(X_{i},x)$ for distribution response, so that the updating function (3.2) can be also re-expressed as:
$\Phi(w_{1})-\Phi(w_{2})=\frac{A(w_{1})-A(w_{2})}{S(w_{1})}+\frac{A(w_{2})\{S(w_{2})-S(w_{1})\}}{S(w_{1})S(w_{2})}.$
Since $A(w_{1})-A(w_{2})=\sum_{i=1}^{n}(W_{i1}-W_{i2})g(X_{i},x)F^{-1}_{i}(z)$,following the Cauchy- Schwartz inequality, we can obtain that:
[…]
where $D_{g,\infty}$ denotes the supremum norm of $g(X_{i},x)$, i.e., $D_{g,\infty}=\max_{i=1,\ldots,n}|g(X_{i},x)|$.
Similarly, we have $\|A(w_{2})\|_{2}\leq nD^{2}_{g,\infty}D_{W}$ and $|S(w_{2})-S(w_{1})|\leq\sqrt{n}D_{g,\infty}\|w_{1}-w_{2}\|_{2}$.
Combining the bounds above, we obtain:
$\|\Phi(w_{1})-\Phi(w_{2})\|_{2}\leq(\frac{n^{\frac{1}{2}}D_{g,\infty}D_{W}}{S_{min}}+\frac{n^{\frac{3}{2}}D^{2}_{g,\infty}D_{W}}{S^{2}_{min}})\|w_{1}-w_{2}\|,$
where $S_{\min}=\min_{w}S(w)$ denotes the minimum possible value.

### v2

@@ Introduction
To assess the performance in the distribution response. For each observation $i=1,\ldots,n$, the covariate $X_{i}$ is independently drawn from a uniform distribution on $[0,1]$. Conditional on $X_{i}$, the true mean parameter $\mu_{i}$ is generated from the normal distribution, $\mu_{i}|X_{i}\sim\mathcal{N}(\mu_{0}+\beta X_{i},v_{1})$. Also, the true standard deviation parameter $\sigma$ is generated from a gamma distribution, $\sigma_{i}|X_{i}\sim{\rm Ga}(\alpha_{i},\lambda_{i})$, where $\alpha_{i}=(\sigma_{0}+\gamma X_{i})^{2}/v_{2}$ is a shape parameter and $\lambda_{i}=v_{2}/(\sigma_{0}+\gamma X_{i})$ is a rate parameter. For a fixed set of quantile levels $\{z_{j}\}$ as an equally spaced sequence starting from 0.1 to 0.9 with an increment of 0.01, expressed as $z_{j}=0.1+0.01\times(j-1)$ for $j=1,2,\ldots,81$. The response variable $Y_{ij}$ at quantile $z_{j}$ is generated via the quantile function of the standard normal distribution as $Y_{ij}=\mu_{i}+\sigma_{i}\Phi^{-1}(z_{j})$, where $\Phi(\cdot)$ denotes the standard normal distribution function. To account for contaminated data, we randomly selected a predetermined number of samples from the observations and introduced a constant shift value to the corresponding response variable at each. Regarding generating contaminated data, we adopted the same strategy as with the matrix response. We randomly select a 10% or 20% dataset and add a 50 or 100 shift in each element of the corresponding distribution observation.
@@
Table 4 summarizes average MISE values averaged over 100 Monte Carlo replications with Monte Carlo standard errors (in parentheses) for the original and robust Fréchet regressions under distribution responses with $n=50$ and $n=100$ across two contamination scenarios. In the absence of contamination, both methods yield comparable MISEs. As the contamination proportion increases, MISEs rise for both methods, but the increase is markedly more pronounced for the original Fréchet regression, especially under large contamination and shifts. For example, at $n=50$ with a contamination proportion of $0.2$ and shift of $100$, the MISE of the robust Fréchet regression is approximately a quarter of that of the original Fréchet regression. Similar to $n=100$, underscoring the superior stability of the robust Fréchet regression in the presence of outliers. The $\lambda$ and $\gamma$ results for the distribution response are also reported in Table 5. Consistent with the matrix-response case, $\gamma$ and $\lambda$ remain large when greater shift value and higher contamination proportions.
@@
Table 5: The value of $\lambda$ and $\gamma$ for distribution response under various data configurations (sample size $n$ and other parameter settings). Values are multiplied by $10^{4}$.
@@
However, our method still has certain limitations. Specifically, it only assesses the overall outliers for each observation as a whole. It does not allow for the evaluation of the outliers of individual components within each observation. For instance, in the case of a matrix response, the anomaly of a single component may lead to the entire matrix being identified as an outlier. Nevertheless, in the estimation process, the presence of other normal components can help decrease the variance of the model. Therefore, extensions addressing this limitation will be considered in our future work.
@@ Appendix A Proof of Proposition 1
under $1\geq W_{i}\geq 0$, the objective function is strictly convex and can be rewritten as $\gamma(W_{i}-\bar{w}_{i})^{2}-\gamma\bar{w}_{i}^{2}+\lambda+\gamma$, where $\bar{w}_{i}=1-(r_{i}(u)-\lambda)/2\gamma$. We consider the case $\bar{w}_{i}\leq 0$, i.e., $r_{i}(u)\geq\lambda+2\gamma$, for which the minimizer achieve at the boundary $W_{i}=0$. We then consider the case $0<\bar{w}_{i}<1$, equivalently, $\lambda<r_{i}(u)\leq\lambda+2\gamma$, where the minimizer is in the interior of the feasible region, and hence $\widetilde{W}_{i}(u)=\bar{w}_{i}$. Last, for the case $\bar{w}_{i}\geq 1$ or equivalently $r_{i}(u)\leq\lambda$, the unconstrained minimum is greater than or equal to 1. The constrained minimizer is the projection of $\bar{w}_{i}$ onto the feasible region, resulting in $\widetilde{W}_{i}(u)=1$.
@@ Appendix B Proof of Proposition 2
$|\widetilde{W}_{i}(u_{1})-\widetilde{W}_{i}(u_{2})|\leq\frac{1}{2\gamma}|r_{i}(u_{1})-r_{i}(u_{2})|\leq\frac{1}{2\gamma}|g(X_{i},x)|L_{d}d(u_{1},u_{2}).$
Hence, it holds that $\|\widetilde{w}(u_{1})-\widetilde{w}(u_{2})\|\leq(2\gamma)^{-1}\sqrt{n}D_{g}L_{d}d(u_{1},u_{2})$. Under (C5), it holds that
@@
where $\rho=\sqrt{n}C_{u}D_{g}L_{d}/2\gamma$. When $\rho<1$, $T$ is a contraction mapping, which completes the proof.
Appendix C Regularity Conditions for Matrix and Distribution Response Models
We check (C4) and (C5) for specific models described in Section 3 and validate the condition, $\rho<1$, given in Proposition 2. To this end, we also assume that $X_{1},\ldots,X_{n}$ are independent and identically distributed random vectors, having mean $\mu_{X}$ and variance-covariance matrix $\Sigma_{X}$. We also assume that $X_{i}$ is sub-Gaussian.
@@ Appendix C Regularity Conditions for Matrix and Distribution Response Models
$\|A(w_{1})-A(w_{2})\|_{F}$ $\leq\|w_{1}-w_{2}\|\Big\{\sum_{i=1}^{n}\|g(X_{i},x)Y_{i}\|_{F}^{2}\Big\}^{1/2}\leq\sqrt{n}\|w_{1}-w_{2}\|D_{g}D_{u},$
where $D_{g}=\max_{i=1,\ldots,n}|g(X_{i},x)|$. We also note that
$\|A(w_{2})\|_{F}\leq\sum_{i=1}^{n}W_{i}|g(X_{i},x)|\cdot\|Y_{i}\|_{F}\leq nD_{g}D_{u},$
and $|S(w_{2})-S(w_{1})|\leq\sqrt{n}D_{g}\|w_{2}-w_{1}\|$ from the Cauchy-Schwartz inequality. To evaluate the behavior of $S(w)$ for some $w\in\mathcal{W}$, we note that $S(w)=\sum_{i=1}^{n}W_{i}+\sum_{i=1}^{n}W_{i}a_{i}$, where $a_{i}=(x-\mu_{X})^{\top}\Sigma_{X}^{-1}(X_{i}-\mu_{X})$ with $E[a_{i}]=0$ is a sub-Gaussian random variable with parameter denoted by $\sigma^{2}(x)$. Since $\sum_{i=1}^{n}W_{i}a_{i}$ is also sub-Gaussian with parameter $\sigma^{2}(x)\sum_{i=1}^{n}W_{i}^{2}$, it follows from (C3) and the Hoeffding inequality that
$P\left(S(w)\geq\xi n-\epsilon\right)\geq 1-2\exp\left(-\frac{\epsilon^{2}}{2\sigma^{2}(x)\sum_{i=1}^{n}W_{i}^{2}}\right)\geq 1-2\exp\left(-\frac{\epsilon^{2}}{2n\sigma^{2}(x)}\right).$
Given that $g_{i}=1+a_{i}$ and the noise term $a_{i}$ is sub-Gaussian with mean zero, the event $g_{i}<0$ corresponds to $a_{i}<-1$, whose probability is exponentially small by standard sub-Gaussian tail bounds. Furthermore, on the event $g_{i}<0$, the $|g_{i}|$ concentrated to zero. Thus, these observations can be asymptotically ignored. Then, with probability approaching to 1, it holds that
$d(\Phi(w_{1}),\Phi(w_{2}))=\|\Phi(w_{1})-\Phi(w_{2})\|_{F}\leq\underbrace{\left(\frac{n^{\frac{1}{2}}D_{g}D_{u}}{\xi n-\epsilon}+\frac{n^{\frac{3}{2}}D^{2}_{g}D_{u}}{(\xi n-\epsilon)^{2}}\right)}_{C_{u}}\|w_{2}-w_{1}\|,$ (8)
@@
Furthermore, it hold that
$\rho=\frac{\sqrt{n}C_{u}L_{d}D_{g}}{2\gamma}=\frac{L_{d}D_{g}^{2}D_{u}}{2\gamma}\left(\frac{n}{\xi n-\epsilon}+\frac{n^{2}D_{g}}{(\xi n-\epsilon)^{2}}\right)\to\frac{L_{d}D_{g}^{2}D_{u}(\xi+D_{g})}{2\gamma\xi^{2}},$
as $n\to\infty$, whereby $\rho<1$ is established when $\gamma$ (tuning parameter in the weight regularization) and $\xi$ (related to the number of non-outliers) are large.
@@
According to the definition of $L_{2}$-Wasserstein distance, we obtain
$|d^{2}(Y_{i},u_{1})-d^{2}(Y_{i},u_{2})|$
$=\bigl(\|u-F^{-1}_{i}(z)\|_{2}-\|u-F^{-1}_{i}(z)\|_{2}\bigr)\bigl(\|u-F^{-1}_{i}(z)\|_{2}+\|u-F^{-1}_{i}(z)\|_{2}\bigr)$
$\leq 2D_{u}d(u_{1},u_{2}),$
whereby (C3) is satisfied with $L_{d}=2D_{u}$. Similarly to Section C.1, we obtain (8) with probability approaching to 1 and $\rho<1$ is satisfied when $\gamma$ and $\delta$ are large.